### v1

@@ Abstract
This work proposes a novel and unified sparse recovery framework, termed the difference of convex Elastic Net (DCEN). This framework effectively balances strong sparsity promotion with solution stability, and is particularly suitable for high-dimensional variable selection involving highly correlated features. Built upon a difference-of-convex (DC) structure, DCEN employs two continuously tunable parameters to unify classical and state-of-the-art models—including Lasso, Elastic Net, Ridge, and $\ell_{1}-\alpha\ell_{2}$—as special cases. Theoretically, sufficient conditions for exact and stable recovery are established under the restricted isometry property (RIP), and a closed-form expression of the DCEN regularization proximal operator is derived. Moreover, two efficient optimization algorithms are developed based on the DC algorithm (DCA) and the alternating direction method of multipliers (ADMM). Within the Kurdyka-Łojasiewicz (KŁ) framework, the global convergence of DCA and its linear convergence rate are rigorously established. Furthermore, DCEN is extended to image reconstruction by incorporating total variation (TV) regularization, yielding the DCEN-TV model, which is efficiently solved via the Split Bregman method. Numerical experiments demonstrate that DCEN consistently outperforms state-of-the-art methods in sparse signal recovery, high-dimensional variable selection under strong collinearity, and Magnetic Resonance Imaging (MRI) image reconstruction, achieving superior recovery accuracy and robustness.
keywords:
sparsity, compressed sensing, nonconvex optimization, DCA, global convergence
{MSCcodes}
90C26, 65K10, 49M29
1 Introduction
@@
In the compressed sensing (CS) framework [candes2006robust, donoho2006compressed], the key challenge in recovering a sparse signal $\bm{x}$ from underdetermined measurements $\bm{y}=\bm{\Phi}\bm{x}$ is choosing a suitable regularization to constrain the solution. Ideally, one minimizes the $\ell_{0}$ norm, i.e., $\min_{\bm{x}}\|\bm{x}\|_{0}$ subject to $\bm{y}-\bm{\Phi}\bm{x}\in\mathcal{E}$, where $\mathcal{E}\subseteq\mathbb{R}^{m}$ models noise (with $\mathcal{E}=\{\mathbf{0}\}$ in the noiseless case). However, $\ell_{0}$ minimization is NP-hard and nonconvex. Candès et al. [candes2006robust] showed that if $\bm{\Phi}$ satisfies the RIP condition, $\ell_{0}$ minimization can be equivalently replaced by the convex $\ell_{1}$ minimization (Basis Pursuit, BP) [chen2001atomic], yielding $\min_{\bm{x}}\|\bm{x}\|_{1}$ subject to $\bm{y}-\bm{\Phi}\bm{x}\in\mathcal{E}$. Yet, $\ell_{1}$ is dominated by large-magnitude components, whereas $\ell_{0}$ is scale-invariant, so the $\ell_{1}$ approximation introduces bias and may cause suboptimal recovery, especially when $\bm{\Phi}$ has highly coherent columns [fannjiang2012coherence].
@@
Recent studies have shown that nonconvex sparsity-promoting measures can not only significantly reduce the required number of measurements, but also achieve superior recovery performance and enhanced sparsity of the solution. Popular nonconvex measures include the $\ell_{p}$ quasi-norm ($p\in(0,1)$) and its variants [candes2008enhancing, chartrand2007exact, daubechies2010iteratively, xu2012l], the minimax concave penalty (MCP) [zhang2010nearly], the capped-$\ell_{1}$ [zhang2010analysis], the smoothly clipped absolute deviation (SCAD) penalty [fan2001variable], the transformed $\ell_{1}$ ($T\ell_{1}$)  [zhang2018minimization], as well as nonconvex measures based on probability density functions [zhou2022unified]. It is worth noting that these nonconvex measures are all separable in structure, i.e., they can be decomposed into a sum of identical one-dimensional functions. While separable nonconvex measures perform well in scenarios with low mutual coherence or low dynamic range, nonseparable nonconvex measures generally exhibit superior sparse recovery performance when the sensing matrix is highly coherent or the signal possesses a high dynamic range [zhou2025recovery].
@@
Among nonseparable measures, ratio-type nonconvex measures have attracted particular attention due to their scale invariance, which enables a closer approximation to the inherently scale-invariant $\ell_{0}$ “norm” (which is not a true norm). Notably, in the one-dimensional case, these two measures coincide. Ratio-type nonconvex measures were first introduced in the work of Hoyer [hoyer2002non], followed by extensive studies on the $\ell_{1}/\ell_{2}$ model [rahimi2019scale] and its associated algorithms [boct2022extrapolated, tao2022minimization, wang2020accelerated, zeng2021analysis], as well as its applications in computed tomography (CT) reconstruction [wang2021limited]. Xu et al. [xu2021analysis] established exact recovery conditions for $\ell_{1}/\ell_{2}$ and provided sufficient conditions for its equivalence to $\ell_{0}$. Other ratio-type measures include $(\ell_{1}/\ell_{2})^{2}$ [jia2025sparse] and the $q$-ratio sparsity measure $(\ell_{1}/\ell_{q})^{\frac{q}{q-1}}(1<q\leq\infty)$ [zhou2021minimization]. A rich body of literature has also investigated the $\ell_{1}/\ell_{q}$ ($q>1$) model, including [huang2018sparse, li2022proximal, wang2022minimizing, wang2023variant]. Despite their advantages in approximating $\ell_{0}$ and promoting sparsity, ratio-type nonconvex measures suffer from highly complex structures and strong nonconvexity. In particular, they are often not globally Lipschitz continuous and may lack coercivity. These issues render theoretical analysis and algorithm design highly challenging. Moreover, some models (e.g., $\ell_{1}/\ell_{2}$) may even tend to produce an abnormally large coefficient while suppressing other nonzero components [rahimi2019scale].
@@
Another important class of nonseparable nonconvex measures is difference-type. The contour of such measures are closer to those of the $\ell_{0}$ norm, making them more effective in promoting sparsity. The $\ell_{1-2}$ (e.g., $\ell_{1}-\ell_{2}$) measure is the most representative example. It was originally proposed in [esser2013method] as a sparsity-inducing penalty for nonnegative least squares problems, and later applied to sparse recovery in [lou2015computing, yin2015minimization], where it demonstrated superior performance over existing measures when the sensing matrix is highly coherent. Subsequently, several other difference-type measures have been proposed and extensively studied, including $\ell_{1}-\alpha\ell_{2}$ [ge2021dantzig, lou2018fast], $\ell_{1}-\beta\ell_{q}$ [huo2023minimization], $\ell_{r}-\alpha\ell_{1}$ [zhou2023rip], $\ell_{q}-\alpha\ell_{p}$ ($p\in(1,2]$[gu2024nonconvex], and weighted $\ell_{q}-\alpha\ell_{p}$ ($q\in(0,1]$, $p\in[q,+\infty]$, $\alpha\in[0,1]$[zhan2025sparse]. Nevertheless, difference-type measures also exhibit inherent limitations. First, as the number of dominant components (ordered by magnitude) in the signal increases, difference-type measures may gradually become biased; for instance, the behavior of $\ell_{1-2}$ tends to approach that of $\ell_{1}$ [ma2017truncated]. Second, classical difference-type measures primarily focus on sparsity itself and often fail to effectively capture the grouping effect when dealing with highly correlated variables, which may result in unstable solutions in practical applications.
@@
To overcome the aforementioned limitations in balancing sparsity promotion and robustness under coherent sensing matrices, this work, inspired by Elastic Net [zou2005regularization], proposes a more general difference-type nonseparable sparsity measure for sparse signal recovery and high-dimensional variable selection. Specifically, we consider the following constrained optimization problem
(1) $\min_{\bm{x}\in\mathbb{R}^{n}}\gamma\bigl(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\bigr)+(1-\gamma)\|\bm{x}\|_{2}^{2}\quad\text{s.t.}\quad\bm{A}\bm{x}=\bm{b},$
@@
(2) $\min_{\bm{x}\in\mathbb{R}^{n}}\frac{1}{2}\|\bm{A}\bm{x}-\bm{b}\|_{2}^{2}+\lambda\left(\gamma\bigl(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\bigr)+(1-\gamma)\|\bm{x}\|_{2}^{2}\right),$
where $\bm{A}\in\mathbb{R}^{m\times n}$ denotes the sensing matrix, $\bm{b}\in\mathbb{R}^{m}$ is the observation vector, and $\bm{x}\in\mathbb{R}^{n}$ is the signal to be recovered. The parameter $\lambda>0$ balances the data fidelity term and the regularization term. Two key tuning parameters are introduced: $\gamma\in(0,1)$ controls the mixing ratio between the $\ell_{1}-\alpha\ell_{2}$ and $\ell_{2}^{2}$ components, while $\alpha\in(0,1)$ adjusts the strength of the nonconvex difference term. The unconstrained regularized problem (2) is referred to as DCEN, and the constrained problem (1) as CDCEN. We primarily investigate the theoretical properties of CDCEN and the optimization methods for solving DCEN. DCEN aims to combine the high-precision recovery capability of nonconvex regularization with the stability of Elastic Net, and seeks an optimal balance between sparsity induction and bias control through adjustable parameters. Notably, DCEN provides a highly unified mathematical framework: by flexibly tuning $\gamma$ and $\alpha$, DCEN reduces to several classical convex and nonconvex regularization models:
@@
In summary, DCEN inherits the advantages of nonconvex measures in sparse approximation while retaining the ability of the Elastic Net to handle correlated variables. Compared with ratio-type measures, DCEN is constructed within the DC framework, which guarantees coercivity of the objective function and enables efficient optimization via DC algorithms (DCA). Figure 1 shows that the objective function of (1), by tuning the parameter $\gamma$, can preserve its sparsity-promoting property while effectively reducing nonconvexity and enhancing smoothness. The main contributions of this work can be summarized as follows:
@@
    Both constrained and unconstrained DCEN models are proposed and some conditions are obtained for ensuring the exact and stable recovery based on RIP, which extends and improvs the existing results in [yin2015minimization, yan2017sparse].
@@
    Global convergence and linear convergence rates of the proposed DCA and ADMM are rigorously established under the KŁ framework, which extend the convergence result of DCA proposed in [yin2015minimization].
@@
    Systematic experiments are provided for noiseless or noisy signal recovery, high-dimensional correlated variable selection, and MRI reconstruction to show that the performance of DCEN consistently is superior to the state-of-the-art methods in terms of both accuracy and stability.
Figure 1: Geometric comparison of $r(\bm{x})=\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}$ and its variant $r_{\gamma}(\bm{x})=\gamma(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2})+(1-\gamma)\|\bm{x}\|_{2}^{2}$.
2 Recovery properties of CDCEN
@@
2.1 Exact and stable recovery
@@
(4) $\|\bm{x}\|_{2}^{2}$ $=\sum_{i=1}^{n}|x_{i}|^{2}=\sum_{i=1}^{n}(x_{\min}+\delta_{i})^{2}=\sum_{i=1}^{n}(x_{\min}^{2}+2x_{\min}\delta_{i}+\delta_{i}^{2})$
$=n\cdot x_{\min}^{2}+2x_{\min}\sum_{i=1}^{n}\delta_{i}+\sum_{i=1}^{n}\delta_{i}^{2}=n\cdot x_{\min}^{2}+2x_{\min}\Delta_{1}+\|\bm{\delta}\|_{2}^{2}.$
@@
(5) $\|\bm{\delta}\|_{2}^{2}=\sum\delta_{i}^{2}\leq\left(\sum\delta_{i}\right)^{2}=\Delta_{1}^{2}.$
@@
(6) $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\geq\|\bm{x}\|_{1}-\alpha(\sqrt{n}x_{\min}+\Delta_{1}).$
@@
(7) $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\geq(n-\alpha\sqrt{n})x_{\min}+(1-\alpha)(\|\bm{x}\|_{1}-nx_{\min}).$
(b) It is worth noting that $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}$ depends only on the nonzero components of $\bm{x}$, that is, $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}=\|\bm{x}_{\Lambda}\|_{1}-\alpha\|\bm{x}_{\Lambda}\|_{2}$. Hence, statement (b) directly follows by applying (a) to $\bm{x}_{\Lambda}$.
Remark 2.3.
The newly derived lower bound is sharper than $(n-\alpha\sqrt{n})x_{\min}$ in [yin2015minimization]. Since $\|\bm{x}\|_{1}-nx_{\min}=\Delta_{1}\geq 0$ and $1-\alpha\geq 0$, the new bound becomes strictly greater than $(n-\alpha\sqrt{n})x_{\min}$ whenever the vector $\bm{x}$ is not flat (i.e., $\|\bm{x}\|_{1}>nx_{\min}$). The proposed bound explicitly exploits the residual term $\Delta_{1}$, which measures the excess of the $\ell_{1}$-norm over its minimum value $nx_{\min}$, thus providing a sharper estimation.
Lemma 2.4.
@@
    $Cx_{\min}\!\left(\alpha n^{\frac{1}{2}-p}\!\left(n^{\frac{1}{2}}-1\right)+(1-\alpha)H_{n,p}\right)\leq\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}$,
@@
    $Cx_{\min}\!\left(\alpha s^{\frac{1}{2}-p}\!\left(s^{\frac{1}{2}}-1\right)+(1-\alpha)H_{s,p}\right)\leq\|\bm{x}_{\Lambda}\|_{1}-\alpha\|\bm{x}_{\Lambda}\|_{2}$,
where $x_{\min}$ is defined as the one in Lemma 2.1 and $H_{n,p}=\sum_{i=1}^{n}i^{-p}$.
Proof 2.5.
@@
(8) $\|\bm{x}\|_{1}=\sum_{i=1}^{n}|x_{i}|\geq\sum_{i=1}^{n}Ci^{-p}x_{\min}=Cx_{\min}\sum_{i=1}^{n}i^{-p}.$
Let $H_{n,p}=\sum_{i=1}^{n}i^{-p}$. Then $\|x\|_{1}\geq Cx_{\min}H_{n,p}$. By Lemma 2.1, substituting (8) into (7) yields $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\geq(\alpha n-\alpha\sqrt{n})x_{\min}+(1-\alpha)Cx_{\min}H_{n,p}$. It follows from $x_{\min}\geq Cn^{-p}x_{\min}$ that $(\alpha n-\alpha\sqrt{n})x_{\min}+\allowbreak(1-\alpha)Cx_{\min}H_{n,p}\geq\allowbreak Cx_{\min}\!\bigl(\alpha n^{\frac{1}{2}-p}(n^{\frac{1}{2}}-1)+\allowbreak(1-\alpha)H_{n,p}\bigr)$.
(b) Statement (b) directly follows by applying (a) to $\bm{x}_{\Lambda}$.
Remark 2.6.
In Lemma 2.4, if the parameters $C$ and $p$ are properly chosen, the new lower bound $Cx_{\min}\left(\alpha s^{\frac{1}{2}-p}\left(s^{\frac{1}{2}}-1\right)+(1-\alpha)H_{s,p}\right)$ becomes tighter than the bound $(s-\alpha\sqrt{s})x_{\min}$. For $C=s^{p}$ with $p\geq 0$, one has $Cx_{\min}\bigl(\alpha s^{\frac{1}{2}-p}(s^{\frac{1}{2}}-1)+\allowbreak(1-\alpha)H_{s,p}\bigr)=\allowbreak x_{\min}\bigl(\alpha(s-\sqrt{s})+\allowbreak(1-\alpha)s^{p}H_{s,p}\bigr)$. Letting $\Delta L=x_{\min}\bigl(\alpha s-\alpha\sqrt{s}+\allowbreak(1-\alpha)s^{p}H_{s,p}\bigr)-\allowbreak(s-\alpha\sqrt{s})x_{\min}$, we have $\Delta L=(\alpha s-s+\allowbreak(1-\alpha)s^{p}H_{s,p})x_{\min}=\allowbreak(1-\alpha)(s^{p}H_{s,p}-s)x_{\min}$. If $0<\alpha<1$ and $p=0$, then $s^{p}H_{s,0}=H_{s,0}=\sum_{i\in\Lambda}\left(\frac{s}{i}\right)^{0}=s$ and $\Delta L=0$. If $0<\alpha<1$ and $p>0$, then for any $i\in\{1,2,\ldots,s-1\}$, it holds that $s/i>1$ and $(s/i)^{p}>1$. Thus, for any $s\geq 2$, one has
(9) $\sum_{i\in\Lambda}\left(\frac{s}{i}\right)^{p}=\left(\frac{s}{1}\right)^{p}+\cdots+\left(\frac{s}{s-1}\right)^{p}+\left(\frac{s}{s}\right)^{p}>s.$
@@
Lemma 2.7.
@@
    (10) $\gamma\big(\alpha n-\alpha\sqrt{n}+(1-\alpha)n^{p}$ $H_{n,p}\big)x_{\min}+(1-\gamma)\|\bm{x}\|_{2}^{2}\leq f_{\alpha,\gamma}(\bm{x})$
    $\leq\gamma(\sqrt{n}-\alpha)\|\bm{x}\|_{2}+(1-\gamma)\|\bm{x}\|_{2}^{2}.$

### v2

@@ Abstract
This work proposes a novel and unified sparse recovery framework, termed the difference of convex Elastic Net (DCEN). This framework effectively balances strong sparsity promotion with solution stability, and is particularly suitable for high-dimensional variable selection involving highly correlated features. Built upon a difference-of-convex (DC) structure, DCEN employs two continuously tunable parameters to unify classical and state-of-the-art models–including LASSO, Elastic Net, Ridge, and $\ell_{1}-\alpha\ell_{2}$–as special cases. Theoretically, sufficient conditions for exact and stable recovery are established under the restricted isometry property (RIP), an oracle inequality and recovery bound are derived for the global solution, and a closed-form expression of the DCEN regularization proximal operator is obtained. Moreover, two efficient optimization algorithms are developed based on the DC algorithm (DCA) and the alternating direction method of multipliers (ADMM). Within the Kurdyka-Łojasiewicz (KŁ) framework, the global convergence of DCA and its linear convergence rate are rigorously established. Furthermore, DCEN is extended to image reconstruction by incorporating total variation (TV) regularization, yielding the DCEN-TV model, which is efficiently solved via the Split Bregman method. Numerical experiments demonstrate that DCEN consistently outperforms state-of-the-art methods in sparse signal recovery, high-dimensional variable selection under strong collinearity, and Magnetic Resonance Imaging (MRI) image reconstruction, achieving superior recovery accuracy and robustness.
I Introduction
@@
In the compressed sensing (CS) framework [4, 10], the key challenge is how to choose a suitable regularization to constrain the solution for recovering a sparse signal $\bm{x}$ from underdetermined measurements $\bm{y}=\bm{\Phi}\bm{x}$. Ideally, one considers the optimization problem $\min_{\bm{x}}\|\bm{x}\|_{0}$ subject to $\bm{y}-\bm{\Phi}\bm{x}\in\mathcal{E}$, where $\mathcal{E}\subseteq\mathbb{R}^{m}$ models noise (with $\mathcal{E}=\{\mathbf{0}\}$ in the noiseless case). However, $\ell_{0}$ minimization is NP-hard and nonconvex. Candès et al. [4] showed that if $\bm{\Phi}$ satisfies the RIP condition, $\ell_{0}$ minimization can be equivalently replaced by the convex $\ell_{1}$ minimization (Basis Pursuit, BP) [8], i.e., $\min_{\bm{x}}\|\bm{x}\|_{1}$ subject to $\bm{y}-\bm{\Phi}\bm{x}\in\mathcal{E}$. Nevertheless, the $\ell_{1}$ approximation leads to both bias and suboptimal recovery because $\ell_{1}$ is dominated by large-magnitude components, especially when $\bm{\Phi}$ has highly coherent columns [13].
@@
Recent studies have shown that nonconvex sparsity-promoting measures not only significantly reduce the required number of measurements, but also achieve superior recovery performance and enhanced sparsity of the solution. Popular nonconvex measures include the $\ell_{p}$ quasi-norm ($p\in(0,1)$) and its variants [5, 7, 9, 39], the minimax concave penalty [44], the capped-$\ell_{1}$ [46], the smoothly clipped absolute deviation penalty [12], and the transformed $\ell_{1}$  [45] as well as nonconvex measures based on probability density functions [48]. It should be noticed that the nonconvex measures mentioned above are all separable in structure, i.e., they can be decomposed into a sum of identical one-dimensional functions. Although the separable nonconvex measures perform well in scenarios with low mutual coherence or low dynamic range, the nonseparable nonconvex measures generally exhibit superior sparse recovery performance when the sensing matrix is highly coherent or the signal possesses a high dynamic range [50].
@@
One important nonseparable measure is the ratio-type nonconvex measure. Such a measure has attracted particular attention due to its scale invariance, which is getting closer as the inherently scale-invariant $\ell_{0}$ “norm” (which is not a true norm). It is worth mentiong that the ratio-type nonconvex measures were first introduced in the work of Hoyer [16], followed by extensive studies on the $\ell_{1}/\ell_{2}$ model [28, 38] and its associated algorithms [2, 31, 35, 42], as well as its applications in computed tomography reconstruction [34]. Other ratio-type nonconvex measures include $(\ell_{1}/\ell_{2})^{2}$ [20] and the $q$-ratio sparsity measure $(\ell_{1}/\ell_{q})^{\frac{q}{q-1}}(1<q\leq\infty)$ [47]. A rich body of literature has also investigated the $\ell_{1}/\ell_{q}$ ($q>1$) model, including [18, 23, 33, 36].Despite having some advantages in approximating $\ell_{0}$ and promoting sparsity, ratio-type nonconvex measures suffer from highly complex structures and strong nonconvexity. Moreover, they are neither globally Lipschitz continuous nor coercive, which render theoretical analysis and algorithm design highly challenging. In addition, some models (e.g., $\ell_{1}/\ell_{2}$) may even tend to produce an abnormally large coefficient while suppressing other nonzero components [28].
@@
Another important nonseparable nonconvex measure is the difference-type. The contour of such a measure is getting closer as the $\ell_{0}$ norm, which makes it more effective in promoting sparsity. The $\ell_{1-2}$ (e.g., $\ell_{1}-\ell_{2}$) measure, originally proposed in [11] as a sparsity-inducing penalty for nonnegative least squares problems and later applied to sparse recovery in [26, 41], has been demonstrated to possess superior performance over existing measures when the sensing matrix is highly coherent. Subsequently, several other difference-type nonconvex measures have been proposed and extensively studied, including $\ell_{1}-\alpha\ell_{2}$ [14, 25], $\ell_{1}-\beta\ell_{q}$ [19], $\ell_{r}-\alpha\ell_{1}$ [49], $\ell_{q}-\alpha\ell_{p}$ ($p\in(1,2]$[15], and weighted $\ell_{q}-\alpha\ell_{p}$ ($q\in(0,1]$, $p\in[q,+\infty]$, $\alpha\in[0,1]$[43]. Nevertheless, difference-type nonconvex measures also exhibit inherent limitations. The first one is that the difference-type nonconvex measures may gradually become biased when the number of dominant components (ordered by magnitude) in the signal increases. For instance, the behavior of $\ell_{1-2}$ tends to approach that of $\ell_{1}$ [27]. The second one is that the difference-type nonconvex measures primarily focus on sparsity itself and often fail to effectively capture the grouping effect when dealing with highly correlated variables, which may result in unstable solutions in practical applications.
@@
To overcome the aforementioned limitations in balancing sparsity promotion and robustness under coherent sensing matrices, in the present paper, inspired by Elastic Net [51], we propose a more general difference-type nonseparable nonconvex sparsity measure for compressed sensing and image reconstruction variable selection. Specifically, we consider the following constrained optimization problem
$\min_{\bm{x}\in\mathbb{R}^{n}}\gamma\bigl(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\bigr)+(1-\gamma)\|\bm{x}\|_{2}^{2}\quad\text{s.t.}\quad\bm{A}\bm{x}=\bm{b},$ (1)
@@
$\min_{\bm{x}\in\mathbb{R}^{n}}\frac{1}{2}\|\bm{A}\bm{x}-\bm{b}\|_{2}^{2}+\lambda\left(\gamma\bigl(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\bigr)+(1-\gamma)\|\bm{x}\|_{2}^{2}\right),$ (2)
where $\bm{A}\in\mathbb{R}^{m\times n}$ denotes the sensing matrix, $\bm{b}\in\mathbb{R}^{m}$ is the observation vector. The parameter $\lambda>0$ balances the data fidelity term and the regularization term. Two key tuning parameters are introduced: $\gamma\in(0,1)$ controls the mixing ratio between the $\ell_{1}-\alpha\ell_{2}$ and $\ell_{2}^{2}$ components, while $\alpha\in(0,1)$ adjusts the strength of the nonconvex difference term. The unconstrained regularized problem (2) is referred to as DCEN, and the constrained problem (1) as CDCEN. DCEN aims to combine the high-precision recovery capability of nonconvex regularization with the stability of Elastic Net, and seeks a balance between sparsity induction and bias control through adjustable parameters. Notably, DCEN provides a highly unified mathematical framework: by flexibly tuning $\gamma$ and $\alpha$, DCEN reduces to several classical convex and nonconvex regularization models:
@@
In summary, DCEN inherits the advantages of nonconvex measures in sparse approximation while retaining the ability of the Elastic Net to handle correlated variables. Compared with ratio-type measures, DCEN is constructed within the DC framework, which guarantees coercivity of the objective function and enables efficient optimization via DC algorithms (DCA). Tuning the parameter $\gamma$ preserves the sparsity-promoting property while effectively mitigating nonconvexity and enhancing smoothness.
Beyond the novelty of the model, the main challenges lie in designing an optimization algorithm for the nonconvex and nonsmooth DCEN with rigorous convergence guarantees and explicit convergence rates, as well as in establishing recovery conditions for CDCEN and DCEN. These challenges necessitate novel technical tools, particularly the careful treatment of the convex–nonconvex hybrid structure jointly governed by $\gamma$ and $\alpha$. In contrast, methods based on a single sparsity measure are generally ill-suited to such hybrid structures. Moreover, our algorithmic design and theoretical analysis differ substantially from existing works [41], whose DC decompositions fail to satisfy the standard DCA assumptions and only guarantee sequence convergence. By contrast, to enhance sparse signal recovery while preserving the grouping effect among variables, we construct a DC decomposition satisfying the standard DCA assumptions and establish both global convergence and explicit convergence rates of the proposed algorithm. The main contributions of this paper can be summarized as follows:
@@
    Both constrained and unconstrained DCEN models are proposed and some conditions are obtained for ensuring the exact and stable recovery based on RIP, which extends and improvs the existing results in [40, 41]. Moreover, we introduce the restricted eigenvalue condition for DCEN and use it to establish the oracle property and recovery bound for the global solution.
@@
    Global convergence and linear convergence rates of the proposed DCA and ADMM are rigorously established under the KŁ framework, which extend the convergence result of DCA proposed in [41].
@@
    Systematic experiments are provided for the noiseless or noisy signal recovery, high-dimensional correlated variable selection, and MRI reconstruction to show that the performance of DCEN consistently is superior to the state-of-the-art methods in terms of both accuracy and stability.
The rest of this work is structured as follows. The next section presents the theoretical analysis of CDCEN. After that in Section IV, a closed-form solution of the proximal operator is given for the DCEN minimization. In Section V, we describe DCA and provide its convergence analysis. In Section VI, we develop an efficient ADMM scheme for solving the DCEN minimization. In Section VII, we report numerical experiment results, before we summarize the results in Section VIII.
II Recovery properties of CDCEN
@@
II-A Exact and stable recovery
@@
$\|\bm{x}\|_{2}^{2}=n\cdot x_{\min}^{2}+2x_{\min}\Delta_{1}+\|\bm{\delta}\|_{2}^{2}.$ (4)
@@
$\|\bm{\delta}\|_{2}^{2}=\sum\delta_{i}^{2}\leq\left(\sum\delta_{i}\right)^{2}=\Delta_{1}^{2}.$ (5)
@@
$\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\geq\|\bm{x}\|_{1}-\alpha(\sqrt{n}x_{\min}+\Delta_{1}).$ (6)
@@
$\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\geq(n-\alpha\sqrt{n})x_{\min}+(1-\alpha)(\|\bm{x}\|_{1}-nx_{\min}).$ (7)
(b) It is worth noting that $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}$ depends only on the nonzero components of $\bm{x}$, that is, $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}=\|\bm{x}_{\Lambda}\|_{1}-\alpha\|\bm{x}_{\Lambda}\|_{2}$. Hence, statement (b) directly follows by applying (a) to $\bm{x}_{\Lambda}$. ∎
Remark 2.
The newly derived lower bound is sharper than $(n-\alpha\sqrt{n})x_{\min}$ in [22]. Since $\|\bm{x}\|_{1}-nx_{\min}=\Delta_{1}\geq 0$ and $1-\alpha>0$, the new bound becomes strictly greater than $(n-\alpha\sqrt{n})x_{\min}$ whenever the vector $\bm{x}$ is not flat (i.e., $\|\bm{x}\|_{1}>nx_{\min}$). The proposed bound explicitly exploits the residual term $\Delta_{1}$, which measures the excess of the $\ell_{1}$-norm over its minimum value $nx_{\min}$, thus providing a sharper estimation.
Lemma 3.
@@
    $Cx_{\min}(\alpha n^{\frac{1}{2}-p}(n^{\frac{1}{2}}-1)+(1-\alpha)H_{n,p})\leq\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}$,
@@
    $Cx_{\min}(\alpha s^{\frac{1}{2}-p}(s^{\frac{1}{2}}-1)+(1-\alpha)H_{s,p})\leq\|\bm{x}_{\Lambda}\|_{1}-\alpha\|\bm{x}_{\Lambda}\|_{2}$,
where $x_{\min}$ is defined as the one in Lemma 1 and $H_{n,p}=\sum_{i=1}^{n}i^{-p}$.
Proof.
@@
$\|\bm{x}\|_{1}=\sum_{i=1}^{n}|x_{i}|\geq\sum_{i=1}^{n}Ci^{-p}x_{\min}=Cx_{\min}\sum_{i=1}^{n}i^{-p}.$ (8)
Let $H_{n,p}=\sum_{i=1}^{n}i^{-p}$. Then $\|x\|_{1}\geq Cx_{\min}H_{n,p}$. Lemma 1 and (8) yield $\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\geq(\alpha n-\alpha\sqrt{n})x_{\min}+(1-\alpha)Cx_{\min}H_{n,p}$. It follows from $x_{\min}\geq Cn^{-p}x_{\min}$ that $(\alpha n-\alpha\sqrt{n})x_{\min}+\allowbreak(1-\alpha)Cx_{\min}H_{n,p}\geq\allowbreak Cx_{\min}\!\bigl(\alpha n^{\frac{1}{2}-p}(n^{\frac{1}{2}}-1)+\allowbreak(1-\alpha)H_{n,p}\bigr)$.
(b) Statement (b) directly follows by applying (a) to $\bm{x}_{\Lambda}$. ∎
Remark 4.
In Lemma 3, if the parameters $C$ and $p$ are properly chosen, the new lower bound $Cx_{\min}\left(\alpha s^{\frac{1}{2}-p}\left(s^{\frac{1}{2}}-1\right)+(1-\alpha)H_{s,p}\right)$ becomes tighter than the bound $(s-\alpha\sqrt{s})x_{\min}$. For $C=s^{p}$ with $p\geq 0$, one has $Cx_{\min}\bigl(\alpha s^{\frac{1}{2}-p}(s^{\frac{1}{2}}-1)+\allowbreak(1-\alpha)H_{s,p}\bigr)=\allowbreak x_{\min}\bigl(\alpha(s-\sqrt{s})+\allowbreak(1-\alpha)s^{p}H_{s,p}\bigr)$. Letting $\Delta L=x_{\min}\bigl(\alpha s-\alpha\sqrt{s}+\allowbreak(1-\alpha)s^{p}H_{s,p}\bigr)-\allowbreak(s-\alpha\sqrt{s})x_{\min}$, we have $\Delta L=(\alpha s-s+\allowbreak(1-\alpha)s^{p}H_{s,p})x_{\min}=\allowbreak(1-\alpha)(s^{p}H_{s,p}-s)x_{\min}$. If $0<\alpha<1$ and $p=0$, then $s^{p}H_{s,0}=H_{s,0}=\sum_{i\in\Lambda}\left(\frac{s}{i}\right)^{0}=s$ and $\Delta L=0$. If $0<\alpha<1$ and $p>0$, then for any $i\in\{1,2,\ldots,s-1\}$, it holds that $s/i>1$ and $(s/i)^{p}>1$. Thus, for any $s\geq 2$, one has
$\sum_{i\in\Lambda}(\frac{s}{i})^{p}=(\frac{s}{1})^{p}+\cdots+(\frac{s}{s-1})^{p}+(\frac{s}{s})^{p}>s.$ (9)
@@
For convenience, we define $\ell(\bm{x})=\frac{1}{2}\|\bm{A}\bm{x}-\bm{b}\|_{2}^{2}$ and $r(\bm{x})=\gamma\big(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\big)+(1-\gamma)\|\bm{x}\|_{2}^{2}$.
Lemma 5.
@@
    $\gamma\big(\alpha n-\alpha\sqrt{n}+(1-\alpha)n^{p}H_{n,p}\big)x_{\min}+(1-\gamma)\|\bm{x}\|_{2}^{2}$ (10)
    $\leq r(\bm{x})\leq\gamma(\sqrt{n}-\alpha)\|\bm{x}\|_{2}+(1-\gamma)\|\bm{x}\|_{2}^{2}.$